OpenStudy (anonymous):

-9.8t2 – 10t + 330=0 how does this work?

OpenStudy (anonymous):

It's the equation for a parabolic motion, where -9.8 is the force of gravity, -10 is initial velocity, and 330 is initial height, in meters. It's the standard position function.

OpenStudy (anonymous):

so if i put in terms of -9.8t^2-10t+330=0....how do i fugure out the (t+-?)(t+-?)

OpenStudy (anonymous):

I'm not sure what you're asking. T is a variable, you need a value to find it out.

OpenStudy (anonymous):

U have to find out the roots of t using b^2\[-(b \pm \sqrt{b ^{2}-4ac})/2a\] which wil give you complex roots in this case .. say u get a+ib , a-ib as the complex roots then you can write the equation as (t-a-ib)*(t-a+ib) ....

OpenStudy (anonymous):

t1=(5\sqrt{3259}-25)/49and t2= (-5sqrt{3259}-25)/49 ]